# Graph Flow Matching: Enhancing Image Generation with Neighbor-Aware Flow Fields

### v2

@@ Appendix D Ablation Study: Diffusion-Reaction with KNN
  • $v_{\mathrm{react}}(X,t)$ is a UNet variant from [44], with code taken from the GitHub repository of [54].
  • $\kappa_{\theta}(X,t)\in\mathbb{R}^{B\times C}$ is a per-example, per-channel diffusivity vector produced by
    $Z=\mathrm{UNet}(X,t),\quad\kappa_{\theta}(X,t)=\sigma\bigl{(}\mathrm{GAP}(Z)\bigr{)}\,,$
@@
  • $L(X)=I-D^{-1/2}A\,D^{-1/2}$ is the normalized graph Laplacian on the batch. We build the adjacency $A\in\mathbb{R}^{B\times B}$ by:
@@
      Computing pairwise cosine similarities $\mathrm{sim}_{ij}=\frac{X_{i}\cdot X_{j}}{\|X_{i}\|\|X_{j}\|}$.
@@
      Masking to retain only these $K$ edges, then row-normalizing so that each row of $A$ sums to one.
  • $X_{\mathrm{flat}}\in\mathbb{R}^{B\times(C\,H\,W)}$ denotes each tensor flattened along spatial dimensions, and $\odot$” broadcasts $\kappa_{\theta}$ across spatial axes.
Equation 10 thus decomposes the velocity into a standard flow-matching “reaction” term $v_{\mathrm{react}}$ and a KNN-based “diffusion” term given by the graph Laplacian $L(X)$ scaled per-channel by $\kappa_{\theta}(X,t)$. This KNN diffusion introduces local coupling across the batch at a cost of $\mathcal{O}(B\,K)$ per step.
To isolate the impact of graph structure from parameter count, we conducted a controlled ablation study on AFHQ-Cat using KNN-based graph diffusion. The architectural hyperparameters for the networks used are provided in Footnote 8. All models were trained for up to 2500 epochs with early-stopping, with a patience of 400 epochs and a learning rate of $10^{-4}$. Table 12 presents these results, comparing GFM variants (with $k=5,10,20$ neighbors) against baseline models without diffusion terms (“noDiff” in the table). Notably, we ensured the baseline models had slightly more parameters ($18.62$M vs. $18.47$M) by increasing their width (number of channels), eliminating the possibility that improved performance stems merely from additional capacity. KNN-based GFM variants consistently outperform these wider baselines across all metrics, demonstrating that neighborhood structure rather than parameter count drives improvement.
@@
Model Batch Total Parameters $v_{\text{diff}}$-net FID ($\downarrow$) FID ($\downarrow$) KID Recall
@@
noDiff (seed $=0$) 100 18.62 0 16.96 17.78 $1.37\cdot 10^{-2}$ 0.21
noDiff (seed $=100$) 64 18.62 0 15.72 15.26 $1.17\cdot 10^{-2}$ 0.32
KNN ($k=5$) 85 18.47 5.54 12.90 12.32 $8.07\cdot 10^{-3}$ 0.38
KNN ($k=10$) 64 18.47 5.54 11.82 11.33 $7.65\cdot 10^{-3}$ 0.38
KNN ($k=20$) 64 18.47 5.54 11.64 10.91 $6.94\cdot 10^{-3}$ 0.36
Table 13: Architectural hyperparameters for the KNN-based diffusion term ablation study (Appendix D) whose results are presented in Table 12. All models use U-Net architectures from [44] [^8] Implementation adapted from https://github.com/annegnx/PnP-Flow and were trained for 2500 epochs with early stopping (patience=400 epochs).
Model Component Base Channels Channel Mult Res Blocks Attn Resolutions
KNN ($k=5,10,20$) $v_{\text{react}}$ $100$ $[1,2]$ $3$ $(32,16)$
$\kappa_{\theta}$-net ($v_{\text{diff}}$) $54$ $[1,2]$ $5$ $(32,16)$
noDiff $v_{\text{react}}$ only $120$ $[1,2]$ $3$ $(32,16)$
Appendix E Timing of Image Generation
The ADM runs were conducted on an Nvidia RTX 4090 (24 GB) GPU and the DiT runs on an Nvidia A6000 (48 GB) GPU. We measured the time required for image generation in Table 14 using the Nvidia A6000 GPU for all models. The number of function evaluations (NFE) and times required to generate 1 image are reported. The corresponding timing results for the checkpoints from [6] were evaluated on our GPU.
@@ Appendix E Timing of Image Generation
Appendix F Comparison with Leading Generative Models
Table 15 compares our GFM variants—ADM+MPNN, ADM+GPS, DiT+MPNN, and DiT+GPS—to a range of leading generative models, including deterministic flow matching (FM) [1], latent diffusion (LDM) [5], wavelet diffusion (WaveDiff) [46], DDPM [17], ImageBART [55], and top GANs (StyleGAN [9], StyleGAN2 [49], ProjectedGAN [56], DiffGAN [57]). As can be observed, augmenting ADM and DiT backbones with our lightweight graph correction term consistently improves FID and recall over their baseline pointwise counterparts. Notably, these ADM+GFM and DiT+GFM models achieve FID and recall on par with the diffusion and GAN pipelines in Table 15, showing that simply adding GFM as a graph based correction yields significant quality gains, without altering the core architecture or training objective.
@@ Appendix F Comparison with Leading Generative Models
Dataset/Model Total Parameters $v_{\text{diff}}$-net FID ($\downarrow$) FID ($\downarrow$) Recall
@@
FM [1] - 00 - $10.54$
LDM [5] - 00 - $4.02$ $0.52$
WaveDiff [46] - 00 - $5.06$ $0.40$
DDPM [17] - 0 - $7.89$
ImageBART [55] - 00 - $7.32$
StyleGAN [9] - 00 - $4.21$
StyleGAN2 [49] - 00 - $3.86$ $0.36$
ProjectedGAN [56] - 00 - $1.59$ $0.44$
@@
LDM [5] - 00 - $4.98$ $0.50$
ImageBART [55] - 00 - $9.57$-
ProjectedGAN [56] - 00 - $3.08$ $0.46$
StyleGAN [9] - 00 - $4.16$ $0.46$
@@
LDM [5] - 00 - $2.95$ $0.48$
DDPM [17] - 00 - $4.90$-
ImageBART [55] - 00 - $5.51$-
ADM [41] - 00 - $1.90$ $0.51$
PGGAN [58] - 00 - $8.34$-
StyleGAN [9] - 00 - $2.35$ $0.48$
ProjectedGAN[56] - 00 - $1.52$ $0.34$
DiffGAN [57] - 00 - $1.43$ $0.58$
Appendix G Qualitative Results

### v3

@@ Appendix C Ablation Study: Diffusion-Reaction with KNN
  • $\mathbf{v}_{\mathrm{react}}(\mathbf{X},t)$ is a UNet variant from (PnPUNet_huang2021variational), with code taken from the GitHub repository of (pnpFlow_martin2025pnpflow).
  • $\kappa_{\theta}(\mathbf{X},t)\in\mathbb{R}^{B\times C}$ is a per‐example, per‐channel diffusivity vector produced by
    $\mathbf{Z}=\mathrm{UNet}(\mathbf{X},t),\quad\kappa_{\theta}(\mathbf{X},t)=\sigma\bigl(\mathrm{GAP}(\mathbf{Z})\bigr)\,$,
@@
  • $\mathbf{L}(\mathbf{X})=\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\,\mathbf{D}^{-1/2}$ is the normalized graph Laplacian on the batch. We build the adjacency $\mathbf{A}\in\mathbb{R}^{B\times B}$ by:
@@
      Computing pairwise cosine similarities $\mathrm{sim}_{ij}=\frac{\mathbf{X}_{i}\cdot\mathbf{X}_{j}}{\|\mathbf{X}_{i}\|\|\mathbf{X}_{j}\|}$.
@@
      Masking to retain only these $K$ edges, then row-normalizing so that each row of $\mathbf{A}$ sums to one.
  • $\mathbf{X}_{\mathrm{flat}}\in\mathbb{R}^{B\times(C\,H\,W)}$ denotes each tensor flattened along spatial dimensions, and $\odot$” broadcasts $\kappa_{\theta}$ across spatial axes.
Equation 9 thus decomposes the velocity into a standard flow‐matching “reaction” term $\mathbf{v}_{\mathrm{react}}$ and a KNN‐based “diffusion” term given by the graph Laplacian $\mathbf{L}(\mathbf{X})$ scaled per‐channel by $\kappa_{\theta}(\mathbf{X},t)$. This KNN diffusion introduces local coupling across the batch at a cost of $\mathcal{O}(B\,K)$ per step.
To isolate the impact of graph structure from parameter count, we conducted a controlled ablation study on AFHQ-Cat using KNN-based graph diffusion. The architectural hyperparameters for the networks used are provided in Table 13. All models were trained for up to 2500 epochs with early stopping, using a patience of 400 epochs (training was terminated if the validation FID did not improve for 400 consecutive epochs) and a learning rate of $10^{-4}$. Table 12 presents these results, comparing GFM variants (with $k=5,10,20$ neighbors) against baseline models without diffusion terms (”noDiff” in the table). Notably, we ensured the baseline models had slightly more parameters (18.6218.62M vs. 18.4718.47M) by increasing their width (number of channels), eliminating the possibility that improved performance stems merely from additional capacity. KNN-based GFM variants consistently outperform these wider baselines across all metrics, demonstrating that neighborhood structure rather than parameter count drives improvement.
@@
noDiff (seed = 0) 100 18.62 0 16.96 17.78 $1.37\times 10^{-2}$ 0.21
noDiff (seed = 100) 64 18.62 0 15.72 15.26 $1.17\times 10^{-2}$ 0.32
KNN ($k=5$) 85 18.47 5.54 12.90 12.32 $8.07\times 10^{-3}$ 0.38
KNN ($k=10$) 64 18.47 5.54 11.82 11.33 $7.65\times 10^{-3}$ 0.38
KNN ($k=20$) 64 18.47 5.54 11.64 10.91 $6.94\times 10^{-3}$ 0.36
@@
Model Component Base Channels Channel Mult Res Blocks Attn. Resolutions
KNN ($k=5,10,20$) $\mathbf{v}_{\text{react}}$ 100 [1,2] 3 (32,16)
$\kappa_{\theta}$-net ($\mathbf{v}_{\text{diff}}$) 54 [1,2] 5 (32,16)
noDiff $\mathbf{v}_{\text{react}}$ only 120 [1,2] 3 (32,16)
Table 13: Architectural hyperparameters for the KNN-based diffusion ablation study (see Appendix C); results in Table 12. Models use the U-Net of (PnPUNet_huang2021variational) and were trained for 2,500 epochs with early stopping (patience = 400).
Appendix D Timing of Image Generation
The ADM runs were conducted on an Nvidia RTX 4090 (24 GB) GPU and the DiT runs on an Nvidia A6000 (48 GB) GPU. We measured the time required for image generation in Table 14 using the Nvidia A6000 GPU for all models. The number of function evaluations (NFE) and times required to generate 1 image are reported. The corresponding timing results for the checkpoints from (dao2023flow) were evaluated on our GPU.
@@ Appendix D Timing of Image Generation
Appendix E Comparison with Leading Generative Models
Table 15 compares our GFM variants—ADM+MPNN, ADM+GPS, DiT+MPNN, and DiT+GPS—to a range of leading generative models, including deterministic flow matching (FM) (lipman2023flowmatchinggenerativemodeling), latent diffusion (LDM) (StableDiffusion_rombach2022high), wavelet diffusion (WaveDiff) (WaveDiff_phung2023wavelet), DDPM (DDPM_ho2020denoising), ImageBART (ImageBART_esser2021imagebart), and top GANs (StyleGAN (FFHQ_dataset_karras2019style), StyleGAN2 (StyleGAN2_karras2020training), ProjectedGAN (ProjectedGAN_sauer2021projected), DiffGAN (DiffGAN_wangdiffusion)). As can be observed, augmenting ADM and DiT backbones with our lightweight graph correction term consistently improves FID and recall over their baseline pointwise counterparts. Notably, these ADM+GFM and DiT+GFM models achieve FID and recall on par with the diffusion and GAN pipelines in Table 15, showing that simply adding GFM as a graph based correction yields significant quality gains, without altering the core architecture or training objective.
Dataset/Model Total Parameters $\mathbf{v}_{\text{diff}}$ FID ($\downarrow$) FID ($\downarrow$) Recall
@@ Appendix E Comparison with Leading Generative Models
FM (lipman2023flowmatchinggenerativemodeling) - 0 - $10.54$ -
LDM (StableDiffusion_rombach2022high) - 0 - $4.02$ $0.52$
WaveDiff (WaveDiff_phung2023wavelet) - 0 - $5.06$ $0.40$
DDPM (DDPM_ho2020denoising) - 0 - $7.89$ -
ImageBART (ImageBART_esser2021imagebart) - 0 - $7.32$ -
StyleGAN (FFHQ_dataset_karras2019style) - 0 - $4.21$ -
StyleGAN2 (StyleGAN2_karras2020training) - 0 - $3.86$ $0.36$
ProjectedGAN (ProjectedGAN_sauer2021projected) - 0 - $1.59$ $0.44$
@@
LDM (StableDiffusion_rombach2022high) - 0 - $4.98$ $0.50$
ImageBART (ImageBART_esser2021imagebart) - 0 - $9.57$ -
ProjectedGAN (ProjectedGAN_sauer2021projected) - 0 - $3.08$ $0.46$
StyleGAN (FFHQ_dataset_karras2019style) - 0 - $4.16$ $0.46$
@@
LDM (StableDiffusion_rombach2022high) - 0 - $2.95$ $0.48$
DDPM (DDPM_ho2020denoising) - 0 - $4.90$ -
ImageBART (ImageBART_esser2021imagebart) - 0 - $5.51$ -
ADM (Dhariwal_ADM_NEURIPS2021_49ad23d1) - 0 - $1.90$ $0.51$
PGGAN (PGGAN_karras2018progressive) - 0 - $8.34$ -
StyleGAN (FFHQ_dataset_karras2019style) - 0 - $2.35$ $0.48$
ProjectedGAN(ProjectedGAN_sauer2021projected) - 0 - $1.52$ $0.34$
DiffGAN (DiffGAN_wangdiffusion) - 0 - $1.43$ $0.58$
[…]
Appendix F Qualitative Results